1 bag of soil = 3 cubic feet garden size l= 9ft ; w=16 ft ; h=3 in how many bags of soil are needed? please give full solution thx
Answer by stanbon(75887)

1 bag of soil = 3 cubic feet garden size l= 9ft ; w=16 ft ; h=3 in how many bags of soil are needed
-----------------
Volume of garden: 9*16*3 cu ft
----
# of bags = (9*16*3)/3 = 9*16 = 144
=======
Cheers,
Stan H.
=====